OpenStudy (anonymous):

what is the solution to he equation? 16+x-11=18-5

OpenStudy (across):

What have you tried?

OpenStudy (anonymous):

welly i kinda dont know were to start

OpenStudy (hanner_b_nanner):

add/subtract like terms.. 16-11=5 18-5=13 5+x=13... -5 -5 x=8

OpenStudy (anonymous):

16+x-11=18-5 16+x-11=13 16+x-11+11=13+11 16+x=24 16-16+x=24-16 x=8

OpenStudy (anonymous):

thx a bunch now i understand how to do it

OpenStudy (anonymous):

You're welcome!